OpenStudy (anonymous):

How many grams of sodium sulfide can be produced when 45.3 g Na react with 105 g S?

OpenStudy (anonymous):

This is what I have so far... This is the balanced equation. The limiting reactant is sodium because it will run out before sodium does. Therefore, sodium is the excess. First, I must how many moles are in the equation. From the balanced equation, I can find that the given amonts equal to a certain amount of moles. I can find that two moles of sodium is equal to 46 g 23g/mol 32.07g/mol 78.05g/mol. 45.3g X 23g/mol = 1.96 mol of Na 1.96 X 1 mol of Na2S/ 2 mol Na = 0.98mol of Na2S 0.98 mol of Na2S X 76.48g/mol where moles cancels out and you get 76.48 g of Na2S

OpenStudy (anonymous):

@fallenangelorchid

OpenStudy (fallenangelorchid):

Write down the balance equation first....In this case i'll be 2Na(s) + S(s) = Na2S(s)

OpenStudy (anonymous):

I have that. Check the link.

OpenStudy (fallenangelorchid):

after that label what's already given to you 2Na(s) + S(s) = Na2S(s) 45.3g 105g m=? 23g/mol 32.07g/mol 78.05g/mol in this case, the limiting reagent is Na and excess is Sulfur because sulfur is available in excess. 45.3g X 23g/mol = 1.96 mol of Na 1.96 X 1 mol of Na2S/ 2 mol Na = 0.98mol of Na2S 0.98 mol of Na2S X 76.48g/mol where moles cancels out and you get 76.48 g of Na2S

OpenStudy (anonymous):

Bro, you just took that from another page, I need a different way, because I don't understand how they done it.

OpenStudy (fallenangelorchid):

Kk then try this : https://answers.yahoo.com/question/index?qid=20101228170118AAU6NkK

OpenStudy (anonymous):

Do you know how to do it or no?

OpenStudy (fallenangelorchid):

@nevermind_justschool Not really bruh

OpenStudy (anonymous):

It's alright, A for effort I guess LoL